Greetings,
I am an older adult who has always struggled with basic algebra. I am not sure that anyone can help me but I would like to ask some questions here in hopes to get a better understanding of math. I do not have any actual math homework; I would just like to learn some math. Here is my first question:
I do not understand how the following algebra problem works:
3/4 + 3/4 = 6/4 I do not understand how the top number can be larger than the bottom number. I also do not understand how to add fractions that have different bottom numbers. I attempted some of the problems on this site but I just got confused.
Thank you in advance for any help you can give me, and I am sorry for asking such silly questions.
Jim

Answer by stanbon(57219)

3/4 + 3/4 = 6/4 I do not understand how the top number can be larger than the bottom number.
---
Think of the 3/4 as a pie that has 4 equal parts.
3/4 means you have three of the 4
---
So (3/4) + (3/4) means you have 3 of the 4 pieces from
a first pie and you have 3 of the 4 pieces from a second
pie.
So you have 6 of those (1/4) pieces .
Thus (3/4) + (3/4) = 6/4
-------
The rule is, " if you have the same denominators in fractions
you are adding, you should just add the numerators and write
that number over the denominator".
So (1/8) + (5/8) = 6/8
Or (2/3) + (3/3) = 5/3
===========================
I also do not understand how to add fractions that have different bottom numbers.
If the denominators are different you have to find a number called
the least common denominator.
---
So to add (2/3) + (1/5) you have to change (2/3) and you have to
change (1/5) so they have the same denominator.
---;
How?
(2/3) = (2/3)(5/5) = 10/15
(1/5) = (1/5)(3/3) = 3/15
----
Now you can add because the denominators are the same:
(2/3) + (1/5) = (10/15)+(3/15) = 13/15
============================================
You can find many, many sites online that will show you
how to add, subtract, multiply, and divide fractions.
Just use Google and search for "adding fractions",
or "multiplying fractions", or any other math you wish
to learn.
================
Cheers,
Stan H.
================